# Shaocong Xu1,2, Pengfei Li2, Xinyu Liu2, Qianpu Sun2, Yang Li2, Shihui Guo1, Zhen Wang3, Bo Jiang3, Rui Wang3, Kehua Sheng3, Bo Zhang3, Hao Zhao2

### v1

@@ A Closer Look
Anomaly Synthesis Pipeline
Table 2: Comparisons with previous methods. C3D refers to the base segmentation model, Cylinder3D (Zhu et al. 2021).
[…]
Figure 6: Comparison with the SOTA using the Risk/AUPR/AUROC/Threashold-Coverage curves.
REAL synthesizes anomaly by resizing the objects presenting in existing scene, as depicted in the upper panel of Fig. 1-(b). However, we observe that this synthesis pipeline fails to adequately represent real anomaly in two aspects: 1) the limited variety of objects in the existing scene makes it challenging to compensate for the lack of semantically rich features through resizing; 2) learning from these synthesized anomalies may lead to a model that classifies real anomaly solely based on point sparsity. As such, we resort to an additional dataset, ShapeNet (Chang et al. 2015), which consists of 220,000 models classified into 3,135 categories, to design our new synthesis pipeline.
Specifically, as shown in Fig. 4, we randomly select and load $G\sim\text{Binomial}(20,0.3)$ objects from ShapeNet into the given scene $\mathbf{x}$, where the probability of not adding any object into $\mathbf{x}$ is also considered. The Binomial(a, b) generates a random value from a binomial distribution with ‘a’ trials and ‘b’ probability of success per trial.
For each object $\mathbf{s}\in\mathbb{R}^{l\times 3}$, we first rotate it into its upright pose. Then, we move $\mathbf{s}$ away from $x^{c}$, which is the center of the scene $\mathbf{x}$, alone the x-axis by $d^{x}\sim\text{Uniform}(r^{\text{min}},0.8*r^{\rm max})$ that $r^{\text{min}}$ is the distance of the closest point from $x^{c}$ and $r^{\text{max}}$ is the furthest point from $x^{c}$. The Uniform(a,b) generates a random value from a uniform distribution with lower bound ‘a’ and upper bound ‘b’. Then, we rotate $\mathbf{s}$ around $x^{c}$ on the xy-plane for $d^{\text{lon}}\sim\text{Uniform}(0,360)$ degrees and denote the resulting object as $\mathbf{s}=(\mathbf{u},\mathbf{v},\mathbf{w})$.
There is a probability that $\mathbf{s}$ does not overlap with $\mathbf{x}$, after moving and rotating. Therefore, if $\mathbf{s}$ is positioned outside of $\mathbf{x}$, the subsequent merge operation is not carried out. Specifically, we stop synthesis process if:
$\begin{gathered}\min\left\{|\bar{u}-i|+|\bar{v}-j|\right\}>\Delta,(i,j,k)\in\mathbf{x}\end{gathered}\vspace{-0.5em}$ (11)
@@
Afterward, since the objects from ShapeNet tend to be smaller in size compared to those in the existing scene, we proceed to resize $\mathbf{s}$ by a factor of $k\sim\text{Uniform}(1,7)$. Following this, we put $\mathbf{s}$ on ground: $\mathbf{\tilde{w}}=\mathbf{w}-\Delta_{w}$, where $\Delta_{w}$ represents the distance between the bottom of $\mathbf{s}$ ($\min\mathbf{w}$) and the point on $\mathbf{x}$ that are the closest to $\mathbf{s}$ when looking from the above. The resulting object is $\mathbf{s}=(\mathbf{u},\mathbf{v},\mathbf{\tilde{w}})$.
Note that terminating the synthesis process at the ‘put on ground’ stage is an alternative design. However, this may lead to trivial solutions in anomaly detection, similar to the issue encountered in REAL, namely, relying on the property of point sparsity for anomaly detection. Hence, we merge $\mathbf{s}$ into $\mathbf{x}$ by adjusting the radii of $\mathbf{x}$, which enables $s$ to adhere to the realistic LiDAR distribution. Specifically, we first represent $\mathbf{s}$ and $\mathbf{x}$ using spherical coordinates, resulting in:
$\begin{gathered}\mathbf{s}:=\{s_{j}=(\text{lon}_{j},\text{lat}_{j},\text{r}_{j})|j=1,\ldots,l\}\\ \mathbf{x}:=\{x_{k}=(\text{lon}_{k},\text{lat}_{k},\text{r}_{k})|k=1,\ldots,n\}\end{gathered}\vspace{-0.5em}$ (12)
@@
Dynamic penalty to further capture the differences between samples. As shown in Fig. 1-(b), the object synthesized through resizing is greater ‘weirdness’ in terms of point sparsity when compared to the ShapeNet object. Hence, to maximize the benefits of the point-wise training paradigm, we introduce a point-wise dynamic penalty loss to handle different samples in a customized manner:
@@
[…]
@@
To ensure a fair comparison, we utilize Cylinder3D (Zhu et al. 2021) as the baseline segmentation model.
Quantitative comparison: As illustrated in Tab. 2, our method achieves a new SOTA in anomalous class segmentation for SemanticKITTI and NuScenes, surpassing the previous SOTA by a significant margin. Specifically, our method achieves an AUPR of 44.68 and an AUROC of 92.69, which exceeded the previous SOTA scores of 8.58 and 7.09 in SemanticKITTI. Moreover, our method achieves an AUPR of 31.58 and an AUROC of 95.24 in NuScenes, which exceeded the previous SOTA scores of 10.38 and 10.74, respectively. However, our method exhibits a degradation of 0.24 in the inlier classification performance, which is lower than the previous method APF (Li and Dong 2023). Nonetheless, we believe that this degradation is trivial compared to the substantial improvement we have achieved in the outlier classification.
Quatitative comparison: Consistent with our quantitative results, the qualitative results, as shown in Fig. 7, demonstrate higher response values in real outlier areas compared to REAL. These findings provide further evidence of the superiority of our point-wise abstaining penalty training paradigm. Since the code for APF is still unavailable, we could not compare our method with it.
Risk-Coverage curve comparison: As shown in Fig. 6-(a), our method achieves a lower risk than REAL when full coverage is applied. With decreasing coverage, both our method and REAL show reduced risk.
Why our method achieve a higher risk compared to REAL when coverage is below 90? (purple dotted line)
There are several reasons behind this phenomenon. Firstly, the threshold for coverage below 90 is close to 0, as shown in Fig. 6-(d). This indicates that our method is confident about all the remaining samples, making it challenging to reject any samples. In such cases, there is a high probability of rejecting true positive samples, which adversely affects the performance. Secondly, as illustrated in Fig. 6-(b) and -(c), the AUPR is 3.711, while the AUROC is approximately 80. These results suggest a significant imbalance between true negative and true positive samples, with the latter often being hard samples with low $\mathbf{p^{o}}$. This imbalance leads to worse risk as the coverage becomes lower than 90.
However, we firmly believe that robust performance in high coverage is more critical than in low coverage since anomalous objects are extremely rare in the real world. Low coverage is characterized by a large number of samples in the perceived scene being classified as anomaly, which is not feasible in real-world applications.
@@
Penalty Dynamic Penalty AUPR AUROC $\text{mIoU}_{\text{old}}$ 43.69 92.51 57.47 44.68 92.69 57.56
Effectiveness of dynamic abstaining penalty. As demonstrated in Tab. 3, the penalty setting is denoted by Equ. 10, while the dynamic penalty setting is represented by Equ. 14. The dynamic penalty setting achieved the best performance. These results provide evidence for the effectiveness of our dynamic penalty supervising signal design.
Table 4: Ablation study on pseudo outlier sample synthesis pipeline in SemanticKITTI.
ShapeNet-based resizing-based AUPR AUROC $\text{mIoU}_{\text{old}}$ 29.14 (14.55 $\downarrow$) 89.56 (2.95 $\downarrow$) 57.31 (0.16 $\downarrow$) 41.82 (1.87 $\downarrow$) 93.04 (0.53 $\uparrow$) 57.30 (0.17 $\downarrow$) 43.69 92.51 57.47
Effectiveness of anomaly synthesis pipeline. We conduct an ablation study of anomaly synthesis pipeline. The results, as presented in Tab. 4, demonstrate a decrease in AUPR by 1.87 and an increase in AUROC by 0.53, when the ShapeNet-based anomaly synthesis pipeline is not utilized. Notably, the AUPR and AUROC values drop significantly by 14.55 and 2.95, respectively, without the resizing-based anomaly synthesis pipeline. This raises the question of whether ShapeNet-based synthesis pipeline is trivial.
To address this concern, we conduct an ablation study of anomaly synthesis pipeline by plotting the risk-coverage curve. As shown in Fig. 5-(right), the blue line (resizing-based) shows an initial decrease in risk followed by a increase. This phenomenon can be further explained as the rejection of true predictions of the inlier samples, resulting in a deterioration of the inlier classification.
On the other hand, the ShapeNet-based synthesis pipeline (orange line) exhibits a consistent reduction in risk as the coverage decreases. This phenomenon can be attributed to the rejection of outlier or misclassified samples. Hence, the ShapeNet-based synthesis pipeline can synthesize objects that are more similar to real anomaly, including those that are prone to misclassification.
Finally, by combining these two synthesis pipeline, our method achieves the best performance with low risk at high coverage and a stable decrease in risk as coverage decreases.
@@ Conclusion
In this work, we first revisit previous methods using the unified theoretical framework of selective classification and propose our new formulation which effectively captures the subtle differences between known and anomalous samples. Furthermore, we design a novel anomaly synthesis pipeline that leverages the richness of the ShapeNet dataset and adheres to the realistic LiDAR distribution to compensate for the lack of semantically rich features. Experimental results demonstrate the superiority of our method across both traditional anomaly detection metrics and newly introduced metrics associated with selective classification. Our method has achieved SOTA performance on public benchmark datasets.

### v3

@@ Introduction
Outlier Synthesis Pipeline
Method SemanticKITTI (Behley et al. 2019) NuScenes (Caesar et al. 2020) AUPR AUROC $\text{mIoU}_{\text{old}}$ AUPR AUROC $\text{mIoU}_{\text{old}}$ Closed-set C3D - - 58.00 - - 58.70 C3D + MSP (Hendrycks and Gimpel 2018) 6.70 74.00 58.00 4.30 76.70 58.70 C3D + MaxLogit (Hendrycks et al. 2019) 7.60 70.50 58.00 8.30 79.40 58.70 C3D + MC-Dropout (Gal and Ghahramani 2016) 7.40 74.70 58.00 14.90 82.60 58.70 C3D + REAL (Cen et al. 2022) 20.08 84.90 57.80 (0.20 $\downarrow$) 21.20 84.50 56.80 C3D + APF (Li and Dong 2023) 36.10 85.60 57.30 (0.70 $\downarrow$) - - - C3D + LiON (ours) 44.68 (8.58 $\uparrow$) 92.69 (7.09 $\uparrow$) 57.56 (0.44 $\downarrow$) 31.58 (10.38 $\uparrow$) 95.24 (10.74 $\uparrow$) 59.11 (0.41 $\uparrow$)
Table 2: Comparisons with previous methods. C3D refers to the base segmentation model, Cylinder3D (Zhu et al. 2021). Comparisons with APF on NuScenes are not conducted, as the results are unavailable and its code is not publicly released.
Figure 5: (a). Comparison with the SOTA using the Risk-Coverage curves; (b). Comparison with different outlier synthesis pipeline using the risk-coverage curve; (c). Relationship between point-wise penalty loss and penalty $\alpha$.
REAL synthesizes outliers by resizing the objects presenting in existing scene, as depicted in the left panel of Fig. 1-(b). However, we observe that this synthesis pipeline fails to represent the real outlier in two aspects: 1) the limited variety of objects in the existing scene makes it challenging to compensate for the lack of semantically-rich information through resizing; 2) learning from these synthesized outliers may lead to a model that classifies real outliers solely based on point sparsity. As such, we resort to an additional dataset, ShapeNet, which consists of 220,000 models classified into 3,135 categories. As shown in Fig. 3, we repeat the synthesis pipeline for $G$ times to insert $G$ outlier objects. For each object, there are six steps as follows:
(a) To synthesize diverse outliers, we first randomly decide the number of outlier objects $G$ according to a Binomial distribution [^4] The Binomial(a, b) generates a random value from a binomial distribution with ‘a’ trials and ‘b’ probability of success per trial.. The specific distribution used is $\text{Binomial}(20,0.3)$. We then load $G$ objects from ShapeNet into the given scene $\mathbf{x}$, where the probability of not adding any object into $\mathbf{x}$ is also considered.
(b) Then, for each object $\mathbf{s}\in\mathbb{R}^{l\times 3}$, we move $\mathbf{s}\in\mathbb{R}^{l\times 3}$ away from scene center $x^{c}$, alone the x-axis by $d^{x}\sim\text{Uniform}(r^{\text{min}},0.8*r^{\rm max})$ [^5] The Uniform(a, b) generates a random value from a uniform distribution with lower bound ‘a’ and upper bound ‘b’. that $r^{\text{min}}$ is the distance of the closest point from $x^{c}$ and $r^{\text{max}}$ is the furthest point from $x^{c}$.
(c) Next, we rotate $\mathbf{s}$ around $x^{c}$ on the xy-plane (around the gravity direction) for $d^{\text{lon}}\sim\text{Uniform}(0,360)$ degrees and denote the resulting object as $\mathbf{s}=(\mathbf{u},\mathbf{v},\mathbf{w})$. There is a probability that $\mathbf{s}$ does not overlap with $\mathbf{x}$, after moving and rotating. Therefore, if $\mathbf{s}$ is positioned outside of $\mathbf{x}$, the subsequent steps are not carried out and we move on the next object. Specifically, we stop synthesis process if:
$\begin{gathered}\min\left\{|\bar{u}-i|+|\bar{v}-j|\right\}>\Delta,(i,j,k)\in\mathbf{x}\end{gathered}$ (11)
@@
(d) Then, since the objects from ShapeNet tend to be smaller in size compared to those in the existing scene, we proceed to resize $\mathbf{s}$ by a factor of $k\sim\text{Uniform}(1,7)$.
(e) Following this, we put $\mathbf{s}$ on ground by setting its last axis to $\mathbf{\tilde{w}}=\mathbf{w}-\Delta_{w}$, where $\Delta_{w}$ represents the distance between the bottom of $\mathbf{s}$ and the point on $\mathbf{x}$ that are the closest to $\mathbf{s}$ along the gravity direction and falls into the x-y plane projection of $s$. The resulting object is $\mathbf{s}=(\mathbf{u},\mathbf{v},\mathbf{\tilde{w}})$.
(f) Finally, to consider the realistic LiDAR’s sampling pattern, we merge $\mathbf{s}$ into $\mathbf{x}$ by adjusting the radii of $\mathbf{x}$. Specifically, we represent $\mathbf{s}$ and $\mathbf{x}$ using spherical coordinates:
$\begin{gathered}\mathbf{s}:=\{s_{j}=(\text{lon}_{j},\text{lat}_{j},\text{r}_{j})|j=1,\ldots,l\}\\ \mathbf{x}:=\{x_{k}=(\text{lon}_{k},\text{lat}_{k},\text{r}_{k})|k=1,\ldots,n\}\end{gathered}$ (12)
@@
Dynamic Penalty
As shown in Fig. 1-(b), the outliers synthesized through resizing are further from the inlier data distribution in terms of point sparsity compared to those synthesized by our pipeline. Hence, to maximize the benefits of the point-wise learning paradigm, we introduce a dynamic penalty loss that handles points in a customized manner:
@@
New evaluation metrics: The loss-based selective risk (Eq. 4) is sub-optimal for serving as an evaluation metric for point-wise classification task because it is not sensitive to the class imbalance which is crucial in this task. Thus, we upgrade it into mIoU-based selective risk, as shown below:
@@
We use Cylinder3D (Zhu et al. 2021) as the baseline segmentation model to ensure a fair comparison. The computational cost remains low with the addition of an outlier detection head, achieving 7 fps on a single NVIDIA 3090 GPU.
Quantitative comparison: As illustrated in Tab. 2, our method achieves a new SOTA in outlier class segmentation for SemanticKITTI and NuScenes, surpassing the previous SOTA by a significant margin. Specifically, our method achieves an AUPR of 44.68 and an AUROC of 92.69, which exceeds the previous SOTA scores by 8.58 and 7.09 in SemanticKITTI. Moreover, our method achieves an AUPR of 31.58 and an AUROC of 95.24 in NuScenes, which exceeds the previous SOTA scores by 10.38 and 10.74, respectively.
Qualitative comparison: Qualitative results, as illustrated in Fig. 6, demonstrate that our method not only locates the outliers more accurately but also does so with greater confidence compared to REAL. Furthermore, in NuScenes (Fig. 6-right ), our method accurately identifies the ‘ego-car’ as outliers, although this category is not incorporated into the training and evaluation phases. These findings provide further evidence of the superiority of our method.
Comparison between LiON and the previous SOTA (APF) from another perspective: In addition to better quantitative results, LiON is easier to implement and requires less computational cost compared to APF. LiON can be implemented by simply adding an extra classifier to an arbitrary LiDAR-based segmentation network and training the network in a single stage using our novel learning paradigm and two different outlier synthesis pipelines. In contrast, APF requires not only an arbitrary segmentation network but also several learnable prototypes, a prototypical constraint module, a generator, a discriminator, and an adversarial mapper, which significantly increase computational costs. Further, APF relies on a two-stage training process, complicating the reproducibility of its results.
Risk-Coverage curve comparison: As shown in Fig. 5-(a), the problem of the CE&CCE, analyzed above, is reflected as a high risk at high coverages, while the CE exhibits an unstable trend when coverage decreases. In contrast, our method achieves a highly competitive risk compared to them at high coverages. With decreasing coverage, our method shows a consistent decline to a plateau in risk.
Why our method achieve a higher risk compared to REAL when coverage is below 90%? (purple dotted line) This is because our method rejects most real outlier samples/points when coverage exceeds 90%. This is proved by that the threshold set to achieve 90% coverage is 0.0051, indicating that all samples with a predicted outlier probability higher than 0.0051 are rejected. This means that when coverage is around 90%, the outlier probabilities of the remaining samples are relatively small, and there are few real outliers left. Therefore, to further decrease the coverage, our method tends to reject more true inliers than true outliers.
However, we believe that robust performance in high coverage is more critical than in low coverage due to the fact that the outlier objects are rare in the real world.
@@
Penalty Dynamic Penalty AUPR $\uparrow$ AUROC $\uparrow$ $\text{mIoU}_{\text{old}}\uparrow$ 43.69 92.51 57.47 44.68 92.69 57.56
@@
Effectiveness of dynamic abstaining penalty. As demonstrated in Tab. 3, the penalty setting is denoted by Eq. 10, while the dynamic penalty setting is represented by Eq. 14. The dynamic penalty setting achieves the best performance. These results provide evidence for the effectiveness of our dynamic penalty design.
ShapeNet Resize AUPR AUROC $\text{mIoU}_{\text{old}}$ 29.14 (14.55 $\downarrow$) 89.56 (2.95 $\downarrow$) 57.31 (0.16 $\downarrow$) 41.82 (1.87 $\downarrow$) 93.04 (0.53 $\uparrow$) 57.30 (0.17 $\downarrow$) 43.69 92.51 57.47
Table 4: Ablation study on outlier synthesis pipeline in SemanticKITTI.
Effectiveness of outlier synthesis pipeline. As shown in Tab. 4, excluding the ShapeNet synthesis pipeline results in a decrease in AUPR by 1.87 and an increase in AUROC by 0.53. Moreover, the AUPR and AUROC drop significantly by 14.55 and 2.95, respectively, without the resize synthesis pipeline. This raises the question of whether ShapeNet synthesis pipeline is trivial.
There are two types of real outliers: those that are distant from the inlier distribution (referred to as far real outliers) and those that lie closer to the inlier distribution (referred to as near real outliers). As shown in Fig. 5-(b), the resize synthesis pipeline shows an initial decrease in risk followed by an increase. This occurs because this pipeline can approximate the far real outlier distribution but struggles with approximating the near real outlier distribution.
On the other hand, our ShapeNet synthesis pipeline exhibits a consistent reduction in risk as the coverage decreases, as it can synthesize outliers that approximate both near and far real outlier distributions. The former result from considering the LiDAR sampling pattern, while the latter stem from considering the diversity of real outliers.
@@ Conclusion
In this work, we first revisit previous methods using the unified lens of selective classification and propose a new formulation which effectively captures the subtle differences between inliers and outliers. Then, we design a novel outlier synthesis pipeline to synthesize diverse and realistic outliers, compensating for the lack of semantically-rich information in point clouds. Experimental results demonstrate the superiority of our method across both traditional outlier detection metrics and newly introduced metrics. Our method has achieved SOTA performance on public benchmark datasets.